OpenStudy (anonymous):

find the equation of the tangent line to the graph of f(x)=ln(1+x^3) at the point (0,0)

OpenStudy (zepp):

What's the derivative of f(x)?

OpenStudy (anonymous):

1 right ?

OpenStudy (zepp):

\[\large \frac{d}{dx}f(x)=\frac{d}{dx}\log(1+x^3)\]

OpenStudy (anonymous):

How do you make fractions? (not the kind with the slanted bar)

OpenStudy (zepp):

\[\large =\frac{d}{dx}f(x)\\\large=\frac{d}{dx}\ln(1+x^3)\\\large= \frac{1}{1+x^3}*f'(1+x^3)\]

OpenStudy (zepp):

Use power rule \[\large= \frac{1}{1+x^3}*f'(1+x^3)\\\large =\frac{3x^2}{1+x^3}\]

OpenStudy (anonymous):

Oh, is it just \frac{a}{b}?

OpenStudy (zepp):

\[\large f'(x)=\frac{3x^2}{1+x^3}\]A tangent line is a linear equation, \(\large y=mx+b\), you want the equation of the line at point (0,0). Plug in 0 in to find the slope; \[\large\frac{3x^2}{1+x^3}=\frac{3(0)^2}{1+(0)^3}=0\]

OpenStudy (zepp):

Now that's easy, just find the equation of the tangent line: \[\large y=mx+b\\\large 0=0x+b\\\large0=0+b\\\large b=0\]

OpenStudy (zepp):

Therefore equation of the tangent line is \(\large y=0\)

OpenStudy (anonymous):

ok I got it.

OpenStudy (zepp):

Sorry, should be \[\large \begin{align} &=\frac{1}{1+x^3}*\frac{d}{dx}(1+x^3)\\ &=\frac{3x^2}{1+x^3} \end{align}\]here: http://puu.sh/K2g3 A little mistake in how I wrote the derivative thing.